# Global Non-Axisymmetric Hall Instabilities in a Rotating Plasma

### v1

@@ 1.1 Basic Equations
The essential fundamental question at present is whether the additional wave modes in a uniform magnetized plasma correspond to additional instabilities in the differentially rotating cylinder. To this end, we attempt to reduce the full linear eigenvalue problem into subsystems that represent the different mode branches of interest. Starting with the whistler modes, we explore a simplified, large-$d_{i}$, Electron-MHD limit of the linearized incompressible Hall-MHD system in section 2. We then proceed to numerical solution of full incompressible Hall-MHD equations (2)–(3) in section 3.
2 Electron-MHD limit
@@
In order to extricate the effect of whistler waves alone, and determine whether they couple to the flow shear to generate instability, we consider a limiting case of the Hall-MHD model. If the ion skin depth, $d_{i}$, is large compared to some characteristic length scale, say $\ell$, then electron motion decouples from the ion motion, and we can neglect the terms containing $\boldsymbol{v}$ ( or $\boldsymbol{\xi}$ ) in the induction equation. This approximation yields the electron-MHD (EMHD) or EMH model (A. V. Gordeev et al., 1994). This approximation can be considered valid in the range of plasma densities such that $d_{i}\gg\ell\gg d_{e}$, or $m_{i}c^{2}/Ze^{2}\gg n\ell^{2}\gg m_{e}c^{2}/Ze^{2}$. The system is quasi-neutral, but the bulk ion motion is unimportant. We maintain the contribution from the inductive ${r\Omega}\hat{\boldsymbol{\phi}}\times\boldsymbol{b}$ in the Ohm’s law, since it is the source of free energy driving instability. Neglecting $\boldsymbol{v}\times\boldsymbol{B}$ while maintaining ${r\Omega}\hat{\boldsymbol{\phi}}\times\boldsymbol{b}$ is consistent for modes where $|\omega|d_{i}\gg V_{A}$. It is shown in sec. 2.2 that for unstable for modes in this system, $|\omega|\sim|\Omega|$, so this approximation is valid at modest $d_{i}$ for weak magnetic fields, i.e. $V_{A}/r\Omega\ll 1$. Addtionally, by neglecting the perturbed bulk plasma motion (i.e. the motion of the ion fluid,) the model describes only the coupling between the sheared, rotating flow and whistler waves.
The EMHD model has the additional advantage of being simpler to analyze than the full HMHD system. In this limit, the radial and axial components of the induction equation become
@@
Here we have defined $|\boldsymbol{k}|^{2}\equiv m^{2}/r^{2}+k^{2}$. The terms in the parentheses of equation 11 include both a term that depends on the flow shear, $r\Omega^{\prime}$, and a term proportional to the curvature of the azimuthal magnetic field, $2B_{\phi}/r$. This equation is not a immediately reducible to one describing a standard special function except when either $m=0$ or $k=0$ with particular choices for the $\Omega(r)$ profile. In particular, using the Keplerian rotation profile, $\Omega\propto r^{-3/2}$, solutions for $\psi$ are not recognizable as standard special functions. Introducing the variable transformation $\psi=u(r)\exp(-1/2r|\boldsymbol{k}|^{2})$ transforms equation (11) to an equation for $u$ in which the coefficient of $u^{\prime}$ vanishes. We also normalize the units using the inner wall dimension $r_{1}$ to normalize the radial coordinate and $1/\Omega_{0}\equiv 1/\Omega(r_{1})$ as a characteristic time scale. Following this procedure, we identify several dimensionless parameters that characterize the problem: $m$, $r_{1}k$, $\mathrm{H}\equiv V_{H}/r_{1}\Omega_{0}=\omega_{A}d_{i}/\Omega_{0}r_{1}$, $\Gamma(r)\equiv r\partial_{r}(\Omega/\Omega_{0})$, and $B_{\phi}/\sqrt{\mu_{0}\rho}\Omega_{0}r_{1}$. The dimensionless quantity $\mathrm{H}\equiv V_{H}/r_{1}\Omega_{0}$ essentially measures the strength of the Hall EMF to the inductive EMF in Ohm’s law. We will reuse $k$ and $\bar{\omega}$ to mean the normalized parallel wave number and Doppler shifted frequency, respectively, i.e. $rk_{1}\rightarrow k$, and $\bar{\omega}/\Omega_{0}\rightarrow\bar{\omega}$. In terms of the normalized quantities, equation (11) is transformed into
@@
where $s\equiv r/r_{1}$ is the normalized radial coordinate, and $\mathrm{H}_{\phi}\equiv(B_{\phi}d_{i}/\sqrt{\mu_{0}\rho})/r_{1}\Omega_{0}$. Since we are dealing with axially periodic cylinders, $k=2n\pi r_{1}/\ell$, for $n\in\mathbb{Z}$. So, for a given $r_{1}$ the parameter $k$ is determined by the axial mode number $n$, and the ratio $\ell/r_{1}$. The outer dimension of the cylinder $r_{2}$ also enters as a parameter via the boundary condition imposed at the outer wall. The aspect ratio, $A\equiv\ell/(r_{2}-r_{1})$, is an additional dimensionless parameter on which the eigenvalues depend. Finally, note that although we have assumed that $d_{i}$ is large in deriving equation (12), the problem still depends on $d_{i}$ implicitly via $\mathrm{H}$. In principle changes in $|\mathrm{H}|$ could result from either a change in $|\boldsymbol{B}|$ or $d_{i}$. However, in order to maintain consistency with the large $d_{i}$ assumption, we must consider that $\mathrm{H}\rightarrow 0\iff\boldsymbol{B}\rightarrow 0$.
Equations (11) and (12) describe the global stability of non-axisymmetric whistler (HSI) instabilities in a differentially rotating cylindrical shear flow threaded with a current-free azimuthal, and uniform vertical magnetic field. In the following, we will consider the local stability predictions of these equations followed by a discussion of global stability criteria and properties of unstable modes. Finally, we proceed with numerical eigenvalue calculations of the EMHD model.
2.1 Local dispersion relation
Although it may sometimes lead to erroneous conclusions (K. E., 1992), an approximate dispersion relation based on a WKB, wave-like ansatz is often used to extract stability information. In some instances the local stability analysis provides a necessary criterion for the existence of unstable modes. The EMHD model is amenable to this approach. Since we have already derived a single ODE that defines the eigenvalue problem, the error introduced by local approximation at this stage is less than would be incurred if the approximation were made earlier. Given a solution to equation (12) that has locally oscillatory behavior, say $u^{\prime\prime}/u=-\delta^{2}$, $\delta$ must satisfy the algebraic relation

### v2

@@ II The Hall-MHD model
One interesting question is whether this taxonomy of waves in a uniform, magnetized plasma can be simply related to the spectrum of instabilities in the differentially rotating cylinder. To this end, in Section III we reduce the full linear eigenvalue problem into subsystems that represent the different mode branches of interest. For the following presentation of numerical results from the Hall-MHD model, it suffices to appreciate that the strength of Hall effect on linear modes is chiefly controlled by the parameter $kd_{i}$.
II.1 Numerical results for Hall-MHD stability
In this section we provide two types of numerical solution of the linearized Hall-MHD equations (7)-(9). The first approach uses a spectral element method to compute the eigenvalues of the incompressible Hall-MHD system in a periodic cylinder. For a given $m,k$ pair, we write $\boldsymbol{v},\boldsymbol{b}\propto\exp(im\phi+ikz-i\omega t)$, and the radial dependence of the solutions is represented using a basis of piecewise differentiable polynomials (typically cubic or quartic). Projecting the equations onto this basis results in a linear eigenvalue problem in terms of the vector of coefficients of the expansion that is solved using the LAPACK ZGGEV routine (Anderson et al., 1999). The problem is formulated so that the resulting eigenvalue is the complex frequency describing the time-dependence of the solution: $\omega=\omega_{r}+i\gamma$. Incompressibility is enforced using an auxiliary variable that is algebraically eliminated in a manner consistent with a penalty method. The magnetic divergence constraint is enforced with a diffusive divergence error correction term. The code we employ is a version of the CYL_SPEC code (Sovinec, 2016) modified to included the Hall term and sheared rotation in the equilibrium.
We compare a subset of the eigenvalue results with linear initial value calculations using the extended MHD code NIMROD (Sovinec et al., 2004). NIMROD (non-ideal magnetohydrodynamics with rotation, open discussion), uses high-order spectral element method to solve both linear and nonlinear extended MHD. The domain is discretized using a 2-D grid of high-order polynomial finite elements in the rzr-z plane, and a pseudospectral collocation method with a Fourier basis for the periodic azimuthal coordinate, $\phi$. The equations are integrated in time using a semi-implicit leapfrog time-stepping scheme. The accuracy and numerical stability of the two-fluid model in NIMROD has been successfully benchmarked with simulations of linear and nonlinear non-ideal MHD instabilities (Sovinec and King, 2010).
In order to compare NIMROD calculations to CYL_SPEC, we compute an effective growth rate from the time-series data output of the code. For a given azimuthal mode number, the velocity and magnetic field are initialized with a small amplitude perturbation. The linear equations are integrated in time until the growth of the most unstable mode can be unambiguously identified. The linear growth rate, $\gamma$, is computed by calculating the change in either the perturbed magnetic and kinetic energies between successive time steps. Concretely, $2\gamma=\log(\Delta\int b^{2}/2\mu_{0}~d^{3}x/\Delta t)$ or $2\gamma=\log(\Delta\int\rho v^{2}~d^{3}x/\Delta t)$, where $b,v$ are the magnitudes of the perturbed magnetic field and perturbed flow, respectively.
We note that the eigenvalue solver does not include electrical resistivity or viscosity. In the NIMROD calculations, however, a small amount of resistivity and viscosity is used for improved numerical stability. The effect of these dissipative terms on the linear growth of the axisymmetric MRI, and non-axisymmetric ‘Magneto-Curvature modes’ modes in MHD is understood (Balbus and Hawley, 1998; Ebrahimi and Haywood, 2025; Haywood and Ebrahimi, 2025). It is also well-established that there can be a competition between the resistive dissipation and the Hall term (Wardle and Salmeron, 2012). The value of magnetic Reynolds number in the NIMROD calculations is $\mathrm{Rm}=2529$, and the magnetic Prandtl number is $\mathrm{Pm}=1$. With these parameters, the effect on the most unstable linear modes is negligible. In the parlance of Wardle and Salmeron (2012), we are firmly in the ‘Hall-MRI’ regime. In this section we focus on the most global (i.e. lowest $m,k$) modes, since these modes would be the least affected by dissipative mechanisms. To assess the effect of varying strength of the Hall term, we compare the results for different values of $d_{i}$ to both the MHD result, and the EMHD limit with $d_{i}=1r_{1}$ for both azimuthal and axial magnetic field configurations in Sections II.2 and II.3, respectively. For the calculations in this section, we normalize lengths by the inner wall radius of the domain, $r_{1}$, and set $r_{2}=5r_{1}$, $\ell=8r_{1}$. Times are normalized by $\Omega_{0}\equiv\Omega(r_{1})$, and velocities by $r_{1}\Omega_{0}$. Since density is uniform, we give magnetic field strengths in terms of the corresponding normalized Alfvén velocities, $V_{A}/r_{1}\Omega_{0}$. This is also referred to as the Lehnert number (Lehnert, 1954)
II.2 Axial Magnetic Field
We first consider the case in which the magnetic field is oriented axially, so $\boldsymbol{B}=B_{z}\hat{\mathbf{z}}$. The linear growth of the axisymmetric, $m=0$ modes in Hall-MHD has been studied (Wardle, 1999; Ebrahimi et al., 2011; Wardle and Salmeron, 2012). Here, we present the linear Hall-MHD results obtained for m=1 modes using the eigenvalue code (CYL_SPEC). We find that it is generally the case that non-axisymmetric modes with $m=1$ are the fastest growing at large magnetic field strengths, but are subdominant to axisymmetric modes at lower magnetic field strengths. The exact relationship depends on both the orientation of the field relative to the direction of rotation, and the value of $d_{i}$. For completeness, we compare the growth rates of the axisymmetric $m=0$ to the non-axisymmetric $m=1$ modes at two values of $d_{i}$ as a function of $V_{A,z}/r_{1}\Omega_{0}$ in figure 3.
For both values of $d_{i}$ the $m=0$ mode has a larger growth rate when $|B_{z}|$ is small, and a higher maximum growth rate. There is an intermediate value of $|B_{z}|$, however, where the $m=0$ mode is stable, and the $m=1$ is the dominant mode. We do not plot the mode frequencies here, since the $m=0$ has $\omega_{r}=0$, and the frequency of the $m=1$ modes are shown in figure 4.
Figure 4 plots the growth rates and frequencies of the fastest growing non-axisymmetric, $m=1$ modes as a function of the magnetic field strength for the MHD model and two values of $d_{i}$ in the Hall-MHD model. The effect of the Hall term is significant even at modest values of $d_{i}$. At $d_{i}=0.1r_{1}$, the ion skin depth is only 2.5% of the radial extent of the disk, and $kd_{i}\approx 0.08$, yet the maximum growth rate of the fastest growing mode is more than twice as large as in the MHD case. The axial field strength at which the maximum growth rate is attained is also larger.
The Hall term destabilizes a different, high-frequency ($\omega\sim\Omega$) branch of non-axisymmetric instability that grows significantly faster than either the $m=1$ MRI or the low frequency global curvature mode in the MHD limit (Ebrahimi and Pharr, 2022).
Figure 4: Growth rates and frequencies of the most unstable mode with $m=1,k=\pi/4$ as a function of the axial magnetic field strength measured by the inverse Alfvénic Mach number. We compare the MHD case in blue to the Hall-MHD result at $d_{i}=0.1$, and $d_{i}=1$ in green and orange, respectively. As the strength of the Hall effect ($d_{i}$) increases, the growth rate develops an asymmetry with respect to the direction magnetic field.
As $d_{i}$ is increased, the dispersion relation develops an increasingly asymmetric dependence on direction of $\boldsymbol{B}$ relative to the axis of rotation as the system transitions from the MHD to the Hall-dominant regime. In the right-handed cylindrical coordinates used, we assume $\boldsymbol{\Omega}\propto\hat{\mathbf{z}}$, so that positive $B_{z}$ corresponds to a magnetic field parallel to the axis of rotation. As $d_{i}$ increases, the $m=1,k=1$ modes with $B_{z}<0$ have lower growth rates than the modes with $B_{z}>0$, but remain unstable at larger values of $|B_{z}|$.
Figure 5 shows the radial profiles of the solutions to the Hall-MHD eigenvalue problem at $d_{i}=0.1$ for $V_{A,z}=0.1r_{1}\Omega_{1}$, $V_{A,z}=0.3r_{1}\Omega_{0}$ and $V_{A,z}=0.5r_{1}\Omega_{0}$.
As the strength of the axial field decreases, the solutions become increasingly oscillatory. This shortening of the radial wavelength is expected from whistler modes. We investigate this point further in section III.
Figure 5: Radial structure of eigenfunctions for Hall-MHD system with $d_{i}=0.1r_{1}$ and axial magnetic fields $V_{A,z}=0.1r_{1}\Omega_{0}$ (left), $V_{A,z}=0.3r_{1}\Omega_{0}$ (center) and $V_{A,z}=0.5r_{1}\Omega_{0}$ (right)
When the axial field is anti-parallel with the axis of rotation, $B_{z}<0$, the whistler modes are not present for $m=1,k=\pi/4$. Instead, the most unstable modes are more similar in radial structure to the low-frequency global ‘curvature modes’ of Ebrahimi and Pharr (2022) that have been modified by the Hall term. As $d_{i}$ increases, the growth rates of the curvature modes for $B_{z}<0$ decrease. Figure 6 shows the eigenmodes from the same three $|V_{A,z}|$ as in figure 5, but with $V_{A,z}<0$.
Figure 6: Radial structure of eigenfunctions for Hall-MHD system with $d_{i}=0.1r_{1}$ and axial magnetic fields $V_{A,z}=-0.1r_{1}\Omega_{0}$ (left), $V_{A,z}=-0.3r_{1}\Omega_{0}$ (center) and $V_{A,z}=-0.5r_{1}\Omega_{0}$ (right)
II.3 Azimuthal Magnetic Field
When the field is purely azimuthal, $B_{\phi}\neq 0,B_{z}=0$, the asymmetric effect of the Hall term is more pronounced. Figure 7 plots the growth rates and frequencies of the fastest growing $m=1,k=\pi/4$ mode as a function of the azimuthal field strength. When $B_{\phi}>0$ we see a significant increase in the growth rates of the most unstable non-axisymmetric modes as $d_{i}$ is increased. Also, with increasing $d_{i}$, the range of $B_{\phi}$ over which the modes are unstable is reduced. Figure 7 only shows $k>0$ modes, but we note that if $k=-\pi/4$, the resulting dispersion relation is mirrored about the vertical axis.
Figure 7: Growth rates and frequencies of the most unstable eigenmode with $m=1,k=\pi/4$ as a function of the azimuthal magnetic field strength at the inner boundary (in units of the inverse Alfvénic Mach number). The MHD result (in blue) and Hall-MHD (HMHD) with of $d_{i}=0.1$ and $d_{i}=1$ are shown.
Based on the behavior of the mode frequencies, we surmise that there are at least two distinct modes present in the calculation. As $|B_{\phi}|\rightarrow 0$, a higher frequency mode with very slow growth dominates.
Figure 8: Radial profiles of the $m=1,k=\pi/4$ HMHD eigenfunctions with $d_{i}=0.1$ with azimuthal fields $V_{A,\phi}=-0.1\Omega_{0}r_{1}$ (left), and $V_{A,\phi}=-0.5\Omega_{0}r_{1}$ (center), and $V_{A,\phi}=-1\Omega_{0}r_{1}$ (right).
As $|B_{\phi}|$ increases, a the lower-frequency mode is destabilized. Unlike the whistler and MRI-type modes, the frequency of this mode is effectively constant as $|B_{\phi}|$ varies.
The polarization of the HMHD ion cyclotron wave in the large $d_{i}$ limit is such that the perturbed flow scales as $|\boldsymbol{v}|\sim|\boldsymbol{k}|d_{i}V_{A}$, and the magnetic field as $|\boldsymbol{b}|\sim|\boldsymbol{B}|$. Thus, using the normalized units, one expects that $|\boldsymbol{v}|/|\boldsymbol{b}|\sim kd_{i}V_{A}/r_{1}\Omega_{0}$. Figure 10 plots the profiles of $v_{r}$ and $b_{r}$ from the most unstable mode at $V_{A,\phi}=-2\Omega_{0}r_{1}$ and $d_{i}=1$. The magnitude of $\boldsymbol{v}$ is approximately double the magnitude of $\boldsymbol{b}$, which is expected for the polarization of the ion-cyclotron for these parameters if $|\boldsymbol{k}|=\sqrt{m^{2}/r_{1}^{2}+k^{2}}$ with $m=1,~k=\pm\pi/4$. Also plotted in figure 10 are the resulting contour plots of the real parts of the perturbed fields after multiplying by $\exp(im\phi)$.
Figure 10: Radial profile (left) and contours in the $r-\phi$ plane (right) of the $m=1,k=\pi/4$ eigenfunction of the Hall-MHD system with $d_{i}=1r_{1}$ at $V_{A,\phi}=-2\Omega_{0}r_{1}$
Table 1: Field strength and growth rates for linear NIMROD calculations in an azimuthal field. Each calculation has $d_{i}=1r_{1}$, $\beta=5\times 10^{6},\mathrm{Rm}\equiv r_{1}^{2}\Omega_{0}/\eta=2529$, $\mathrm{Pm}=1$. For each value of $V_{A,\phi}/r_{1}\Omega_{0}$, we compare to the results for the $k=\pm\pi/4$ modes from CYL_SPEC
$V_{A,\phi}/r_{1}\Omega_{0}$ NIMROD (initial value) CYL_SPEC (eigenvalue)
$\gamma/\Omega_{0}$ $|k|$ (observed) $\gamma/\Omega_{0}$ $k$ (input)
-2.0 0.03 $\pi/4$ 0.03 $\pi/4$
-0.5 0.05 $\pi/4$ 0.05 $\pi/4$
-0.1 0.13 $\pi/2$ 0.07 $\pi/4$
0.1 0.13 $\pi/2$ 0.07 $\pi/4$
0.5 0.05 $\pi/4$ 0.05 $\pi/4$
2.0 0.03 $\pi/4$ 0.03 $\pi/4$
Table 1 summarizes the linear growth rates of the $m=1$ mode from NIMROD calculations in at a few values of azimuthal field strength for $d_{i}=1r_{1}$. The growth rates agree with the eigenvalue results at $V_{A,\phi}/r_{1}\Omega_{0}=\pm 0.5$ and $V_{A,\phi}/r_{1}\Omega_{0}=\pm 2$, but NIMROD reports higher growth for modes at $V_{A,\phi}/r_{1}\Omega_{0}=\pm 0.1$ magnetic field strength. This discrepancy is unsurprising, and owes to the fact that the axial dimension in NIMROD is discretized using finite elements. For a given azimuthal mode number, the calculated solution is a superposition of eigenmodes with several axial wavenumbers. The calculation of the growth rate is dominated by the axial wavenumber with the largest growth rate, which, for low magnetic fields, increases with $k$. This point is illustrated in figure 11, where we show contours of the radial component of $\boldsymbol{v}$ in the rzr-z plane from NIMROD calculations at several field strengths. Visual inspection of the NIMROD solution at $V_{A,\phi}/r_{1}\Omega_{0}=0.1$ indicates that the dominant mode has axial wavenumber $k=2\pi/4$. Since we have only computed growth rates from CYL_SPEC at $k=\pm\pi/4$, we should not expect agreement with the larger $k$ solution from NIMROD. Table 1 also shows both the value of $|k|$ estimated by counting the number of vertical nodes in the perturbed NIMROD solution, and the input $k$ for the corresponding CYL_SPEC calculation.
Figure 11: Contours of perturbed radial flow, $v_{r}$, in the rzr-z plane from linear NIMROD calculations at $V_{A,\phi}/r_{1}\Omega_{0}=0.1,0.5$, and 22 (left to right). The lowest magnetic field strength case shows a faster growing mode with $k=\pi/2$, while the higher field cases show $m=1,k=-\pi/4$.
At larger magnetic field strength, however, the NIMROD calculation finds that the fastest growing mode is the same $m=1,k=\pm\pi/4$ that was calculated using CYL_SPEC. Figure 12 plots the radial component of the $m=1$ velocity and magnetic field perturbations from the NIMROD calculation with $V_{A,\phi}/r_{1}\Omega_{0}=0.5$ and $V_{A,\phi}/r_{1}\Omega_{0}=2$.
Figure 12: Contours of $v_{r}$ and $b_{r}$ in the $r-\phi$ plane from NIMROD calculations with $d_{i}=1.0$ when $V_{A,\phi}=0.1r_{1}\Omega$ (left) and $V_{A,\phi}=2r_{1}\Omega_{0}$ (right). The lower field case is more localized near the inner boundary. The higher field case agrees with the contours from CYL_SPEC in figure 10
The faster growing, high-$k$ mode observed when $V_{A,\phi}/r_{1}\Omega_{0}=1$ is radially localized near the point of maximum flow shear in the domain. At the larger field strength, the global mode is the most unstable.
III Electron-MHD limit – unstable whistler modes
@@
In order to isolate the effect of whistler waves alone, and determine whether they couple to the flow shear to generate instability, we consider a limiting case of the Hall-MHD model. If the ion skin depth, $d_{i}$, is large compared to the wave-length corresponding to the dynamics of interest, say $\lambda$, then electron motion decouples from the ion motion, and we can neglect the terms containing $\boldsymbol{v}$ ( or $\boldsymbol{\xi}$ ) in the induction equation. This approximation yields the electron-MHD (EMHD) or EMH model (Gordeev et al., 1994). This model can be considered valid in the range of plasma densities such that $d_{i}\gg\lambda\gg d_{e}$, or $m_{i}c^{2}/Ze^{2}\gg n\lambda^{2}\gg m_{e}c^{2}/Ze^{2}$. In terms of the parameters in equations (10)-(13), it is appropriate to equate $\lambda$ with either the azimuthal or axial wavelength of the perturbations. Thus, the EMHD model should accurately describe perturbations with $d_{i}m/r,~kd_{i}\gg 1$. In this limit, the system remains quasi-neutral, but the bulk ion motion is unimportant for the dynamics. Applying these assumptions to our problem of interest, we maintain the contribution from the inductive ${r\Omega}\hat{\boldsymbol{\phi}}\times\boldsymbol{b}$ in the Ohm’s law, since it is the source of free energy driving instability. Neglecting $\boldsymbol{v}\times\boldsymbol{B}$ while maintaining ${r\Omega}\hat{\boldsymbol{\phi}}\times\boldsymbol{b}$ is consistent for modes where $|\omega|d_{i}\gg V_{A}$. It is shown in Section III.2 that for unstable for modes in this system, $|\omega|\sim|\Omega|$, so this approximation is valid at modest $d_{i}$ for weak magnetic fields, i.e. $V_{A}/r\Omega\ll 1$. Additionally, by neglecting the perturbed bulk plasma motion (i.e. the motion of the ion fluid,) the model describes only the coupling between the sheared, rotating flow and whistler waves. Applying these assumptions, equation (8) becomes
$\partial_{t}\boldsymbol{b}+\Omega\partial_{\phi}\boldsymbol{b}-r\Omega^{\prime}\hat{\boldsymbol{\phi}}b_{r}=\frac{d_{i}}{\sqrt{\mu_{0}\rho}}\left(\nabla\times\boldsymbol{b}\cdot\nabla\boldsymbol{B}-\boldsymbol{B}\cdot\nabla\nabla\times\boldsymbol{b}\right).$ (14)
Equation (14), together with $\nabla\cdot\boldsymbol{B}=0$, defines the EMHD approximation we will consider here.
In terms of the variables $\psi,\varphi,b_{z}$ used in Section II, the radial and axial components of the equation (14) become
@@
Here we have defined $|\boldsymbol{k}|^{2}\equiv m^{2}/r^{2}+k^{2}$. The terms in the parentheses of equation 18 include both a term that depends on the flow shear, $r\Omega^{\prime}$, and a term proportional to the curvature of the azimuthal magnetic field, $2B_{\phi}/r$. This equation is not immediately reducible to one describing a standard special function except when either $m=0$ or $k=0$ with particular choices for the $\Omega(r)$ profile. In particular, using the Keplerian rotation profile, $\Omega\propto r^{-3/2}$, solutions for $\psi$ are not immediately recognizable as standard special functions. Introducing the variable transformation $\psi=u(r)\exp(-1/2r|\boldsymbol{k}|^{2})$ transforms equation (18) to an equation for $u$ in which the coefficient of $u^{\prime}$ vanishes. We also normalize the units using the inner wall dimension $r_{1}$ to normalize the radial coordinate and $1/\Omega_{0}\equiv 1/\Omega(r_{1})$ as a characteristic time scale. Following this procedure, we identify several dimensionless parameters that characterize the problem: $m$, $r_{1}k$, $\mathrm{H}\equiv V_{H}/r_{1}\Omega_{0}=\omega_{A}d_{i}/\Omega_{0}r_{1}$, $\Gamma(r)\equiv r\partial_{r}(\Omega/\Omega_{0})$, and $B_{\phi}/\sqrt{\mu_{0}\rho}\Omega_{0}r_{1}$. The dimensionless quantity $\mathrm{H}$ essentially measures the strength of the Hall EMF to the inductive EMF generated by flow shear in Ohm’s law. The latter rotates the radial component of the perturbed magnetic field towards the azimuthal direction, while the former produces right-handed rotations of both components. It is the net effect of these two terms that results in overstable whistler modes.
We will reuse $k$ and $\bar{\omega}$ to mean the normalized parallel wave number and Doppler shifted frequency, respectively, i.e. $r_{1}k_{1}\rightarrow k$, and $\bar{\omega}/\Omega_{0}\rightarrow\bar{\omega}$. In terms of the normalized quantities, equation (18) is transformed into
@@
where $s\equiv r/r_{1}$ is the normalized radial coordinate, and $\mathrm{H}_{\phi}\equiv(B_{\phi}d_{i}/\sqrt{\mu_{0}\rho})/r_{1}^{2}\Omega_{0}$. Since we are dealing with axially periodic cylinders, $k=2n\pi r_{1}/\ell$, for $n\in\mathbb{Z}$. So, for a given $r_{1}$ the parameter $k$ is determined by the axial mode number $n$, and the ratio $\ell/r_{1}$. The outer dimension of the cylinder $r_{2}$ also enters as a parameter via the boundary condition imposed at the outer wall. The aspect ratio, $A\equiv\ell/(r_{2}-r_{1})$, is an additional dimensionless parameter on which the eigenvalues depend. Finally, note that although we have assumed that $d_{i}$ is large in deriving equation (19), the problem still depends on $d_{i}$ implicitly via $\mathrm{H}$. In principle changes in $|\mathrm{H}|$ could result from either a change in $|\boldsymbol{B}|$ or $d_{i}$. However, in order to maintain consistency with the large $d_{i}$ assumption, we must consider that $\mathrm{H}\rightarrow 0\iff\boldsymbol{B}\rightarrow 0$.
Equations (18) and (19) describe the global stability of non-axisymmetric whistler (HSI) instabilities in a differentially rotating cylindrical shear flow threaded with a current-free azimuthal, and uniform axial magnetic field. In the following, we will consider the local stability predictions of these equations followed by a discussion of global stability criteria and properties of unstable modes. Finally, we proceed with numerical eigenvalue calculations of the EMHD model using the CYL_SPEC code.
III.1 Local dispersion relation
Although it may sometimes lead to erroneous conclusions (Knobloch, 1992), an approximate dispersion relation based on a WKB, wave-like ansatz is often used to extract stability information. In some instances the local stability analysis provides a necessary criterion for the existence of unstable modes. The EMHD model is amenable to this approach. Since we have already derived a single ODE that defines the eigenvalue problem, the error introduced by local approximation at this stage is less than would be incurred if the approximation were made earlier. Given a solution to equation (19) that has locally oscillatory behavior, say $u^{\prime\prime}/u=-\delta^{2}$, $\delta$ must satisfy the algebraic relation